# From “Worse is Better” to Better: Lessons from a Mixed Methods Study of Ansible’s Challenges

### v1

@@ Abstract.
Infrastructure as Code (IaC) tools have transformed the way IT infrastructure is automated and managed, but their growing adoption has also exposed numerous challenges for practitioners. In this paper, we investigate these challenges through the lens of Ansible, a popular IaC tool. Using a mixed methods approach, we investigate challenges, obstacles, and issues faced by practitioners. We analyze 59,157 posts from Stack Overflow, Reddit, and the Ansible Forum to identify common pain points, complemented by 16 semi-structured interviews with practitioners of varying expertise levels.
@@
copyright: none
@@ 1. Introduction
Despite the significant advantages of IaC, practitioners face several challenges. Since IaC involves writing code, errors can be introduced, leading to misconfigurations that affect system security (Rahman et al., 2019, 2021; Reis et al., 2022; Saavedra and Ferreira, 2022; Saavedra et al., 2023; Opdebeeck et al., 2023) and system reliability  (GitLab, 2014; Stack Exchange Network, 2014; gooeyblob, 2016; Honeycomb, 2021; Jan David Nose, 2023; Wikimedia, 2017). For example, due to bugs in their IaC scripts, GitHub experienced an outage of their DNS infrastructure (Fryman, 2014) and Amazon Web Services lost around 150 million dollars after issues with their S3 billing system (Hersher, 2017). To address this, there has been an effort by the research community to categorize, identify and repair defects (Saavedra and Ferreira, 2022; Saavedra et al., 2023; Rahman et al., 2021; Reis et al., 2022; Hassan et al., 2024; Begoug et al., 2024; Sotiropoulos et al., 2020; Opdebeeck et al., 2022; Rahman et al., 2023; Lepiller et al., 2021; Schwarz et al., 2018; Sharma et al., 2016; Bessghaier et al., 2024; Opdebeeck et al., 2023; Zerouali et al., 2023; Rahman and Williams, 2018; Rahman et al., 2019, 2020; Saavedra et al., 2025). In addition, IaC tools themselves present challenges, such as unintuitive debugging mechanisms, inconsistent behaviors, and complex learning curves. Prior research has explored challenges faced by IaC practitioners, identifying common issues and limitations of existing tools (Rahman et al., 2018; Guerriero et al., 2019; Begoug et al., 2023; Tanzil et al., 2023). However, existing studies often take a broad perspective, examining challenges across multiple DevOps or IaC tools, which can dilute insights specific to any one tool. Studies focused on a single IaC tool have not, so far, included Ansible.
In this paper, we aim to deepen the understanding of the challenges faced by Ansible users. Ansible is currently the most widely used configuration management tool (Stack Overflow, 2024) and, while it is known that Ansible has issues, such as difficulties with variable scoping (Opdebeeck, 2024) and debugging (Hacker News User gizdan, 2021), the specifics of these challenges and their impact on practitioners remain underexplored. Using a mixed methods approach, we analyze a large corpus of online discussions and conduct semi-structured interviews to systematically examine the obstacles faced by Ansible users. Our study is the first to provide a detailed account of these issues and to offer actionable recommendations for improvement.
@@
  • Insights from 16 semi-structured interviews with Ansible practitioners of varying expertise levels, providing qualitative perspectives on usability and technical challenges.
@@ 2. Ansible
Inventory. In Ansible, the hosts to be configured are listed in an inventory, typically stored as an INI or YAML file. In this file, hosts can be grouped by assigning them specific tags. An example of such an inventory is shown in Figure 1.
@@
Variables and Templates. Ansible uses variables to manage system differences, allowing the same variable name to hold different values for each host. These variables can be used in playbooks to adjust task behavior based on the target host. They can be defined in various ways, such as within playbooks or in Ansible’s inventory, as shown in Figure 1. Additionally, Ansible automatically gathers special variables called facts, which provide system information such as operating system details. Variable precedence is determined by the location where the variable is defined and follows a hierarchy of 22 rules (Ansible project contributors, 2025). Ansible leverages Jinja2 templating to enable dynamic expressions and access to variables and facts (Ansible project contributors, 2025), both within the playbooks and in configuration files used via the template module.
@@
Push-based model. Unlike other configuration management tools, such as Chef and Puppet, Ansible follows a push-based configuration model. Typically, configuration management tools use a pull-based model that involves a server that stores configurations and target hosts that retrieve them. However, in a push model, the server actively sends configurations to the target hosts rather than waiting for them to request updates. Since target hosts do not initiate communication, push-based systems such as Ansible do not require an agent on each machine. Instead, a server, commanded by the Ansible user and with SSH access to all hosts, deploys Ansible modules to bring each host to the desired state.
@@
Example. Figure 2 presents excerpts of an Ansible playbook for configuring a web server. This playbook consists of a single play. Here, we show three tasks and one handler. In line 2, the playbook specifies that it applies to all target hosts listed in the inventory. Line 2 defines the use of the webserver role, ensuring that all tasks within this role are executed. Line 2 demonstrates the use of an Ansible fact to display the operating system of the target host, using Jinja2 syntax for variable access, if the condition specified in line 2 is met. In line 2, the playbook specifies that if the state of the nginx package changes, the handler Restart Nginx (line 2) should be triggered. Line 2 employs a Jinja template to dynamically generate a default web page, allowing the use of Ansible variables for customization.
{minted}
[fontsize=]text [all:vars] ansible_user=admin
[web_servers] web1.example.com ansible_host=192.168.1.10 http_port=80 nginx_version=1.18 web2.example.com ansible_host=192.168.1.11 http_port=8080 nginx_version=1.18
Figure 1. Example of an Ansible INI inventory.
{minted}
[linenos=true,fontsize=,xleftmargin=2em, escapeinside=!!]yaml — - name: Set up and configure web server hosts: all !! gather_facts: yes roles: [’webserver’] # Role to set up a web server !! tasks: - name: Display system facts debug: msg: ”The system OS is ansible_facts[’os_family’] .” !! when: ansible_facts[’os_family’] == ’Debian’ !!
- name: Ensure Nginx is installed apt: name: nginx notify: Restart Nginx !!
- name: Deploy an index.html template template: !! src: templates/index.html.j2 dest: /var/www/html/index.html notify: Reload Nginx
(…) handlers: - name: Restart Nginx !! service: name: nginx state: restarted (…)
@@ 3. Mixed Methods Study
In this section, we outline our research questions and describe the methodology used to answer them. Our study follows a mixed methods approach, summarized in Figure 3. First, we collected posts from Q&A platforms and used topic modeling to understand the challenges, obstacles, and issues faced by Ansible practitioners. Then, we conducted semi-structured interviews to gain deeper insights into participants’ experiences with Ansible.
@@ 3.2.1. Data Collection
We start by collecting posts about challenges, obstacles, and issues faced by Ansible practitioners from three Q&A platforms: StackOverflow (Stack Exchange, 2025b), Reddit (Reddit, 2025), and the Ansible Forum (Red Hat, 2025a) ( 1 in Figure 3). To collect posts from StackOverflow, we used the Stack Exchange API (Stack Exchange, 2025a) and retrieved all posts with the tag Ansible, which corresponds to a total of 23,000 posts. For Reddit, we used the PullPush API (PullPush team, 2025) to collect all 7,187 posts on the subreddit r/ansible that contained the keyword Ansible in addition to keywords possibly related to issues, such as error, help, or slow. For the Ansible Forum, we used web scraping techniques to collect 28,970 posts in the Get Help and Archives categories. We included the category Archives in our analysis, as it contains all posts from the Ansible Google Group, the predecessor of the Ansible Forum, which were later migrated to the Ansible Forum (Red Hat, 2025b). We filtered out posts related to announcements and releases by the Ansible team. In total, we collected 59,157 posts from the three Q&A platforms, forming our final dataset. Data collection took place between January 24, 2025, and February 1, 2025.
@@ 3.2.3. Topic Refinement
TopicGPT generates a large number of topics that, while mostly relevant, are too granular for practical analysis. Moreover, Pham et al. (Pham et al., 2024) identified instances of topics that are duplicated or out of scope. TopicGPT performs an automated topic refinement step to overcome these issues. However, in our first experiments, we achieved unsatisfactory results, even when using more powerful LLMs such as o3-mini-2025-01-31. For example, in one of our test runs, this approach was unable to merge the following pair of topics with a cosine similarity of $0.96000.9600$, which TopicGPT uses to measure semantic proximity:
@@
Next, we applied an iterative manual refinement process similar to the merging step used in TopicGPT. In each iteration, we systematically evaluated topic pairs, starting with those exhibiting the highest cosine similarity and proceeding to those with the lowest similarity down to a predefined threshold for that iteration. Similarly to TopicGPT, we used Sentence-Transformer embeddings (Reimers and Gurevych, 2019) to calculate cosine similarity. Then, for each pair, one author determined whether the topics should be merged and, if so, whether they should be combined into one of the existing topics or assigned to a newly created topic. The criterion for merging topics was the presence of significant semantic overlap between them. If a topic was merged in a given iteration, pairs that included the merged topic were excluded from further evaluation within that iteration. A second author independently reviewed the merged topics, documenting any disagreements or ambiguous cases requiring further discussion. After this independent review, both authors met to resolve discrepancies and reach a consensus. The process was repeated six times until no further merges were identified, resulting in a candidate topic list. In the first three iterations, we applied a threshold of 0.7, while in the last three iterations, we lowered the threshold to 0.6. After the merging process, we reduced our topic list from 153 to 105 topics.
@@
An author reviewed each topic, determining whether it should be removed, renamed, or modified. Subsequently, a second author independently reviewed the proposed changes, documenting any disagreements or ambiguous cases requiring further discussion. After this independent review, both authors met to resolve discrepancies and reach a consensus. This process resulted in the final list of topics composed of 87 topics ( 5 in Figure 3).
@@ 3.2.5. Validation
Two authors independently performed confirmation labeling on a sample of 300 (post, topic) pairs ( 7 in Figure 3). The inter-rater agreement, measured using Cohen’s Kappa (Cohen, 1960), was approximately 0.29, which according to Landis and Koch’s interpretation (Landis and Koch, 1977), is classified as a fair agreement. A third author performed confirmation labeling on the (post, topic) pairs where the initial two authors’ judgments diverged. The final judgment was decided by majority vote. Ultimately, TopicGPT’s topic assignment achieved a precision ( 8 in Figure 3) of approximately 82.67% (see Supplemental Materials for full topic list).
@@ 3.2.6. Analysis
To draw relevant insights across our topic analysis and the interviews, we merged the final list of topics derived from TopicGPT with the thematic codes identified in the interview data (described in Section 3.3). Specifically, we integrated the topics with a set of Challenges codes by aligning each topic with the most closely related theme. This merging process was iterative and followed a double-blind procedure: initially, two coders independently grouped the topics into themes. Finally, both coders met to discuss disagreements and reach a consensus. In analyzing the merged data, we examined the relative frequency of each topic across the resulting themes and investigated how these frequencies evolved. In our analysis, we restricted our scope to the last five years of Q&A data to better capture current and emerging concerns, totaling 23,215 recent posts. In addition to recent posts, we leveraged the data older than five years to identify evolution trends.
@@ 3.3. Interviews
In addition to the large-scale analysis of Q&A posts, we conducted semi-structured interviews to gain a deeper understanding of practitioners’ day-to-day experiences with Ansible. Below, we describe the recruitment process, structure of the interviews, and analysis. All interview materials are in the Supplemental Materials.
@@ 3.3.1. Participants
We categorize participants based their level of expertise with Ansible. For this we use years of experience due to support in previous work (LaToza and Myers, 2010; Roehm et al., 2012) and also to mitigate the Dunning-Kruger Effect, where users overestimate their skills in self-assessments (Mahmood, 2016). We define three levels of expertise with Ansible: Beginner (less than two years of experience, n = 4); Intermediate (two to five years of experience, inclusive, n = 5); and Experienced (more than five years of experience, n = 7).
@@
We recruited 16 participants through a pre-screening survey distributed via email and various social media platforms, including LinkedIn, X (formerly Twitter), Discord, and Slack ( 1 in Figure 3). To meet our eligibility criteria, participants had to have prior experience using Ansible. Following this stage, participants selected their preferred interview modality: a text-based or video-based session conducted via Zoom. The availability of both formats was intended to address known challenges in recruiting software developers for research studies (Baltes and Diehl, 2016), a challenge we also encountered in our study. A total of 118 out of 120 survey participants were eligible and, among these, 16 were available for an interview (an interview rate of 14%). Most of the participants (11 out of 16) opted for a text interview. Several participants mentioned that they would only be able to participate in the interview if they could do it through text. As an incentive, the interviewed participants could enter a raffle for a \$100 gift card.
The supplementary material contains full details on the demographics of participants. In summary, participants reported anywhere from 1 to more than 9 years of Ansible experience, with many currently using Ansible (12). Most identified as male (15), ranged in age from 18–34 (9), and held at least a bachelor’s degree (13). Almost all had a job in CS and had a formal CS education (15).
@@ 3.3.2. Interviews
We conducted semi-structured interviews to obtain a nuanced understanding of the participants’ experiences with Ansible ( 2 in Figure 3). Semi-structured interviews are the most widely used interviewing method (DiCicco-Bloom and Crabtree, 2006) and allow adaptability to explore emerging themes and seek clarifications (Lazar et al., 2017). We piloted the protocol with four participants.Voice-based interviews lasted around 30 minutes. All interviews were recorded and automatically transcribed with participant consent using Zoom’s built-in transcription functionality. Research team members manually reviewed and corrected the resulting transcripts. Each interview followed a structured protocol that encompassed five key thematic areas: contextual background, perceived advantages, challenges, recommendations, and future of Ansible. The interview started with general background questions We then moved the discussion towards the advantages of Ansible and adoption challenges. Then, we focused the discussion on Ansible’s limitations. The interview concluded with exploring recommendations to improve Ansible and future perspectives. We actively sought clarification whenever responses were ambiguous or required elaboration.

### v2

# The Ultimate Configuration Management Tool? Lessons from a Mixed Methods Study of Ansible’s Challenges

@@ Abstract.
Infrastructure as Code (IaC) tools have transformed the way IT infrastructure is automated and managed, but their growing adoption has also exposed numerous challenges for practitioners. In this paper, we investigate these challenges through the lens of Ansible, a popular IaC tool. Using a mixed methods approach, we investigate challenges, obstacles, and issues faced by practitioners. We analyze 59,157 posts from Stack Overflow, Reddit, and the Ansible Forum to identify common pain points, complemented by 20 semi-structured interviews with practitioners of varying expertise levels.
@@
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06copyright: none
@@ 1. Introduction
Despite the significant advantages of IaC, practitioners face several challenges. Since IaC involves writing code, errors can be introduced, leading to misconfigurations that affect system security (Rahman et al., 2019, 2021; Reis et al., 2022; Saavedra and Ferreira, 2022; Saavedra et al., 2023; Opdebeeck et al., 2023) and system reliability  (GitLab, 2014; Stack Exchange Network, 2014; gooeyblob, 2016; Honeycomb, 2021; Jan David Nose, 2023; Wikimedia, 2017). For example, due to bugs in their IaC scripts, GitHub experienced an outage of its DNS infrastructure (Fryman, 2014) and Amazon Web Services lost around 150 million dollars after issues with its S3 billing system (Hersher, 2017). To address this, there has been an effort by the research community to categorize, identify and repair defects (Saavedra and Ferreira, 2022; Saavedra et al., 2023; Rahman et al., 2021; Reis et al., 2022; Hassan et al., 2024; Begoug et al., 2024; Sotiropoulos et al., 2020; Opdebeeck et al., 2022; Rahman et al., 2023; Lepiller et al., 2021; Schwarz et al., 2018; Sharma et al., 2016; Bessghaier et al., 2024; Opdebeeck et al., 2023; Zerouali et al., 2023; Rahman and Williams, 2018; Rahman et al., 2019, 2020; Saavedra et al., 2025). In addition, IaC tools themselves present challenges, such as unintuitive debugging mechanisms, inconsistent behaviors, and complex learning curves. Prior research has explored challenges faced by IaC practitioners, identifying common issues and limitations of existing tools (Rahman et al., 2018; Guerriero et al., 2019; Begoug et al., 2023; Tanzil et al., 2023). However, existing studies often take a broad perspective, examining challenges across multiple DevOps or IaC tools, which can dilute insights specific to any one tool. Studies focused on a single IaC tool have not, so far, included Ansible.
In this paper, we aim to deepen the understanding of the challenges faced by Ansible users. Ansible is currently the most widely used configuration management tool (Stack Overflow, 2024) and, although Ansible is known to have issues, such as difficulties with variable scoping (Opdebeeck, 2024) and debugging (Hacker News User gizdan, 2021), the specifics of these challenges and their impact on practitioners remain underexplored. Using a mixed methods approach, we analyze a large corpus of online discussions and conduct semi-structured interviews to systematically examine the obstacles faced by Ansible users. Our study is the first to provide a detailed account of these issues and to offer actionable recommendations for improvement.
@@
  • Insights from 20 semi-structured interviews with Ansible practitioners of varying expertise levels, providing qualitative perspectives on usability and technical challenges.
@@ 2. Ansible
Variables and Templates. Ansible uses variables to manage system differences, allowing the same variable name to hold different values for each host. These variables can be used in playbooks to adjust task behavior based on the target host. They can be defined in various ways, such as within playbooks or in Ansible’s inventory. Additionally, Ansible automatically gathers special variables called facts, which provide system information such as operating system details. Variable precedence is determined by the location where the variable is defined and follows a hierarchy of 22 rules (Ansible project contributors, 2025). Ansible leverages Jinja2 templating to enable dynamic expressions and access to variables and facts (Ansible project contributors, 2025), both within the playbooks and in configuration files used via the template module.
@@
Push-based model. Unlike configuration management tools such as Chef and Puppet, Ansible follows a push-based configuration model. Typically, configuration management tools use a pull-based model that involves a server that stores configurations and target hosts that retrieve them. However, in a push model, the server actively sends configurations to the target hosts rather than waiting for them to request updates. Since target hosts do not initiate communication, push-based systems such as Ansible do not require an agent on each machine. Instead, a server, commanded by the Ansible user and with SSH access to all hosts, deploys Ansible modules to bring each host to the desired state.
@@
Example. Figure 1 presents excerpts of an Ansible playbook for configuring a web server. This playbook consists of a single play. Here, we show three tasks and one handler. In line LABEL:line:hosts, the playbook specifies that it applies to all target hosts listed in the inventory. Line LABEL:line:roles defines the use of the webserver role, ensuring that all tasks within this role are executed. Line LABEL:line:fact demonstrates the use of an Ansible fact to display the operating system of the target host, using Jinja2 syntax for variable access, if the condition specified in line LABEL:line:condition is met. In line LABEL:line:notify, the playbook specifies that if the state of the nginx package changes, the handler Restart Nginx (line LABEL:line:restartnginx) should be triggered. Line LABEL:line:template employs a Jinja template to dynamically generate a default web page, allowing the use of Ansible variables for customization.
1
2---
3- name: Set up and configure web server
[…]
5gather_facts: yes
[…]
7tasks:
8- name: Display system facts
9debug:
[…]
13apt:
14name: nginx
[…]
18src: templates/index.html.j2
19dest: /var/www/html/index.html
20notify: Reload Nginx
[…]
22handlers:
[…]
24service:
25name: nginx
26state: restarted
27(...)
@@ 3. Mixed Methods Study
In this section, we outline our research questions and describe the methodology used to answer them. Our study follows a mixed methods approach (see Figure 2). First, we collected posts from Q&A platforms and used topic modeling to understand the challenges faced by Ansible practitioners. We then conducted semi-structured interviews to gain deeper insights into the participants’ experiences.
@@ 3.2.1. Data Collection
We start by collecting posts about challenges, obstacles, and issues faced by Ansible practitioners from three Q&A platforms: StackOverflow (Stack Exchange, 2025b), Reddit (Reddit, 2025), and the Ansible Forum (Red Hat, 2025b) ( 1 in Figure 2). To collect posts from StackOverflow, we used the Stack Exchange API (Stack Exchange, 2025a) and retrieved all posts with the tag Ansible, which corresponds to a total of 23,000 posts. For Reddit, we used the PullPush API (PullPush team, 2025) to collect all 7,187 posts on the subreddit r/ansible that contained the keyword Ansible in addition to keywords possibly related to issues, such as error, help, or slow. For the Ansible Forum, we used web scraping techniques to collect 28,970 posts in the Get Help and Archives categories. We included the category Archives in our analysis, as it contains all posts from the Ansible Google Group, the predecessor of the Ansible Forum, which were later migrated to the Ansible Forum (Red Hat, 2025c). We filtered out posts related to announcements and releases by the Ansible team. In total, we collected 59,157 posts from the three Q&A platforms, forming our final dataset. Data collection took place between January 24, 2025, and February 1, 2025.
@@ 3.2.3. Topic Refinement
TopicGPT produces numerous topics that, while generally relevant, are often overly granular, duplicated, or out of scope (Pham et al., 2024). To mitigate these limitations, it incorporates an automated topic refinement step. However, in our first experiments, we achieved unsatisfactory results, even when using more powerful LLMs such as o3-mini-2025-01-31. For example, in one of our test runs, this approach was unable to merge the following pair of topics with a cosine similarity of $0.9600$, which TopicGPT uses to measure semantic proximity:
@@
Next, we applied an iterative manual refinement process similar to the merging step used in TopicGPT. In each iteration, we systematically evaluated topic pairs, starting with those exhibiting the highest cosine similarity and proceeding to those with the lowest similarity down to a predefined threshold for that iteration. Similarly to TopicGPT, we used Sentence-Transformer embeddings (Reimers and Gurevych, 2019) to calculate cosine similarity. Then, for each pair, one author determined whether the topics should be merged and, if so, whether they should be combined into one of the existing topics or assigned to a newly created topic. The criterion for merging topics was the presence of significant semantic overlap between them. If a topic was merged in a given iteration, pairs that included the merged topic were excluded from further evaluation within that iteration. A second author independently reviewed the merged topics, documenting any disagreements or ambiguous cases requiring further discussion. After this independent review, both authors met to resolve discrepancies and reach a consensus. The process was repeated six times until no further merges were identified, resulting in a candidate topic list. In the first three iterations, we applied a cosine similarity threshold of 0.7, while in the last three iterations, we lowered the threshold to 0.6. After the merging process, we reduced our topic list from 153 to 105 topics.
@@
An author reviewed each topic, determining whether it should be removed, renamed, or modified. Subsequently, a second author independently reviewed the proposed changes, documenting any disagreements or ambiguous cases requiring further discussion. After this independent review, both authors met to resolve discrepancies and reach a consensus. This resulted in a final list of 87 topics ( 5 in Figure 2, see Supplemental Materials for the full topic list).
@@ 3.2.5. Validation
Two authors independently performed confirmation labeling on a sample of 300 (post, topic) pairs ( 7 in Figure 2). The inter-rater agreement, measured using Cohen’s Kappa (Cohen, 1960), was approximately 0.29, which according to Landis and Koch’s interpretation (Landis and Koch, 1977), is classified as a fair agreement. A third author performed confirmation labeling on the (post, topic) pairs where the initial two authors’ judgments diverged. The final judgment was decided by majority vote. Ultimately, TopicGPT’s topic assignment achieved a precision ( 8 in Figure 2) of approximately 82.67%.
@@ 3.2.6. Analysis
To draw relevant insights across our topic analysis and the interviews, we merged the final list of topics derived from TopicGPT with the thematic codes identified in the interview data (Section 3.3). Specifically, we integrated the topics with a set of Challenges codes by aligning each topic with the most closely related theme. This process was iterative and followed a double-blind procedure: first, two coders independently grouped the topics into themes. Then, both coders met to discuss disagreements and reach a consensus. In analyzing the merged data, we examined the relative frequency of each topic across the resulting themes and investigated how these frequencies evolved. We restricted our scope to the last five years of Q&A data to better capture current and emerging concerns, totaling 23,215 posts. In addition to recent posts, we leveraged the data older than five years to identify evolution trends.
@@ 3.3. Interviews
In addition to the large-scale analysis of Q&A posts, we conducted semi-structured interviews to gain a deeper understanding of practitioners’ day-to-day experiences with Ansible. In the following section, we describe the recruitment process, the structure of the interviews, and the analysis. All interview materials are in the Supplemental Materials.
@@ 3.3.1. Participants
We categorize participants based on their level of expertise with Ansible. For this we use years of experience due to support in previous work (LaToza and Myers, 2010; Roehm et al., 2012) and also to mitigate the Dunning-Kruger Effect, where users overestimate their skills in self-assessments (Mahmood, 2016). We define three levels of expertise with Ansible: Beginner (less than two years of experience, n = 5); Intermediate (two to five years of experience, inclusive, n = 7); and Experienced (more than five years of experience, n = 8).
@@
We recruited 20 participants through a pre-screening survey distributed via email and various social media platforms, including LinkedIn, X (formerly Twitter), Discord, and Slack ( 1 in Figure 2). To meet our eligibility criteria, participants had to have prior experience using Ansible. Following this stage, participants selected their preferred interview modality: a text-based or video-based session conducted via Zoom. The availability of both formats was intended to address known challenges in recruiting software developers for research studies (Baltes and Diehl, 2016), a challenge we also encountered in our study. A total of 118 out of 127 survey participants were eligible and, among these, 20 were available for an interview (an interview rate of 15.7%). Most of the participants (12 out of 20) opted for a text interview. Several participants mentioned that they would only be able to participate in the interview if they could do it through text. As an incentive, the interviewed participants could enter a raffle for a \$100 gift card.
The supplementary material contains full details on the demographics of the participants. In summary, participants reported anywhere from 1 to more than 9 years of Ansible experience, with many currently using Ansible (15). Most identified as male (17), ranged in age from 18–34 (11), and held at least a bachelor’s degree (17). Almost all had a job in CS and had a formal CS education (17).
@@ 3.3.2. Interviews
We conducted semi-structured interviews to obtain a nuanced understanding of the participants’ experiences with Ansible ( 2 in Figure 2). Semi-structured interviews are the most widely used interviewing method (DiCicco-Bloom and Crabtree, 2006) and allow adaptability to explore emerging themes and seek clarifications (Lazar et al., 2017). We piloted the protocol with four participants. Both text and voice-based interviews required around 30 minutes of attention time. The latter were recorded and automatically transcribed with participant consent using Zoom’s built-in transcription functionality. Research team members manually reviewed and corrected the resulting transcripts. Each interview followed a structured protocol that encompassed five key thematic areas: contextual background, perceived advantages, challenges, recommendations, and future of Ansible. The interview started with general background questions. We then moved the discussion towards the advantages of Ansible and adoption challenges. Then, we focused the discussion on Ansible’s limitations. The interview concluded with exploring recommendations to improve Ansible and future perspectives. We actively sought clarification whenever responses were ambiguous or required elaboration, both in textual and voice-based interviews.